# Spin Squeezing Enhanced Quantum Magnetometry with Nitrogen-Vacancy Center Qutrits

### v1

@@ I Introduction
In this study, we explore the potential enhancements in quantum metrological performance achievable by transitioning from two-level (qubits) to three-level (qutrits) spin systems and incorporating spin squeezing within typical quantum magnetometry settings, specifically NV center Ramsey interferometry. We adopt a foundational approach, comparing qubit versus qutrit Ramsey quantum magnetometry scenarios, both with and without spin squeezing, in terms of Quantum Fisher Information (QFI). Considering the influence of quantum dephasing and relaxation during the spin system’s evolution, our findings suggest that qutrits, coupled with the additional step of spin squeezing, can outperform qubits in terms of QFI, provided that the spin squeezing magnitude, the timing of squeezing and free evolution are optimally determined.
@@ II Model
In a scenario where there exists an unknown, potentially weak magnetic field within a defined region, our objective is to measure this magnetic field with the utmost precision possible. To accomplish this, we utilize one or two NV centers as our probes. The NV center, characterized as an optically active color defect, results from a substitutional nitrogen impurity and a neighboring carbon vacancy in the diamond lattice. Comprising six electrons from nitrogen and surrounding carbon atoms, the negatively charged NV center is described as a spin-1 system. We opt to explore the full capabilities of the NV center by modeling it as a qutrit system. This choice is motivated by the potential advantages that qutrits offer in metrology when compared to qubits [29, 30]. Its ground state is a spin triplet (${}^{3}A_{2}$) denoted as |$\ket{Sm_{S}}$ with $S=1$ and $m_{S}=0,\pm 1$. The NV center Hamiltonian captures these properties, while the excited-state triplet (${}^{3}E$) is acknowledged at a higher energy level and is not explored in this context [31, 32].
@@
Figure 1: Illustration of Ramsey-type sequence with and without introduction of spin squeezing. Note that the times mentioned in the figure correspond to those considered in our study.
@@ II.1 Free Hamiltonian and Initial states
When incorporating spin squeezing (details provided in the following subsections), the initial state, |$\ket{\psi_{0}}$, is first driven through spin squeezing, and then subjected to the $\pi/2$ pulse, as illustrated in Fig. 1.
@@ II.2 Spin Squeezing Generation
where $c_{1}$ is the squeezing constant. The Hamiltonian, as depicted in Eq. (6), is referred as a one-axis twisting Hamiltonian. The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin. This, combined with the optimal ground state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
@@
Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$.
@@ II.3 Dynamics
The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins. These nuclear spins introduce a dynamic and unpredictable backdrop, contributing to the environmental noise that impacts the precision of measurements performed on the NV center. We assume very low temperatures relative to the NV center transition frequencies. Dephasing and thermalization are incorporated through the dissipative rate. They are respectively given as follows,
@@ III.2 Spin Squeezing Parameter
with $I_{3}$ being the three dimensional unit matrix. To calculate spin squeezing, we calculate the mean spin squeezing direction $\hat{n}$ for the state and then take the minimal variance in the perpendicular direction of $\hat{n}$,
@@
The presence of spin squeezing is characterized by $\xi^{2}<1$. Spin squeezing is closely tied to the creation of entanglement [37, 49, 50]. Additionally, a previous study suggests that initial squeezing can aid in entanglement [51]. Our results are consistent with this statement (Appendix A).
Figure 3: The curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
@@ IV Results
Initially, we investigate the typical Ramsey sequence for one and two NV centers, both starting in the initial state (Eq. 5). Subsequently, after a $\pi/2$ pulse, we allow the system to undergo evolution through the free Hamiltonian $\hat{H}_{0}$ for a duration of $T_{\mathrm{fr}}=20$ $\mu$s, $T_{\mathrm{fr}}$ being the free evolution time. We choose this time taking into account the dissipative time that falls within the order of tenth of microseconds. To further enhance precision, starting in the initial state (Eq. 5), we consider introducing spin-squeezing before the $\pi/2$ pulse and the free evolution. The two scenarios are illustrated in Fig. 1.
@@ IV.1 Analysis with spin-1 formalism
[…]
Fig. 2 compares two NV centers under different conditions: one without spin squeezing and the other with spin squeezing applied at varying strengths, specifically with squeezing constants of $c_{1}=0.25$ GHz and $c_{1}=2.5$ GHz. In the absence of noise, the QFI is larger when spin squeezing is applied, and it increases further with a higher squeezing constant, $c_{1}$ (Fig. 2a). When environmental noise is introduced with $\gamma_{\text{d}}=0.1\gamma_{\text{t}}$, spin squeezing remains beneficial, particularly at a lower squeezing constant ( $c_{1}=0.25$ GHz) (Fig. 2b). Applying moderate squeezing, sufficiently large to achieve its effects, shows advantages by being more robust against the impact of environmental noise. However, within the first 4 $\mu$s, a higher squeezing constant provides more benefit. As the dephasing rate increases to $\gamma_{\text{d}}=\gamma_{\text{t}}$, the advantage of squeezing is lost (Fig. 2c).
The effect of spin squeezing on the NV centers is further illustrated in Fig. 3. The squeezing parameter, initially around 0.5 (below 1) at the beginning of free evolution, increases due to dephasing and thermalization during this period. Appendix A demonstrates that spin squeezing also leads to entanglement. The duration of the application of spin squeezing, denoted by $T_{\mathrm{sq}}$, corresponds to the time when the squeezing parameter reaches its minimum value.
The initial state is oriented along the $z$-axis. When spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis. A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid. The effect of the magnetic field $B$ is to precess the spin around the $z$-axis, and the $x$-axis twisting deforms the noise cone, effectively narrowing the spin’s error ellipse. This sequence is illustrated in Fig. 4.
@@
Experimentally, we typically measure populations or ($\hat{S}_{z}$) values. However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the Symmetric Logarithmic Derivative (SLD), measuring $\hat{S}_{z}$ could potentially be suboptimal.
While spin squeezing enhance magnetic sensing capabilities, they are also highly sensitive to environmental noise, particularly dephasing. Analyzing the behavior of the QFI curves under such noise conditions is crucial. By identifying the time at which the QFI reaches its maximum, we can optimize experimental design and ensure measurements are performed at the peak sensitivity.
Figure 5: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$. The NV centers system are described by spin-1/2 formalism.
@@ IV.2 Analysis with spin-1/2 formalism
As in the spin-1 formalism, if spin squeezing is added, it is applied before the $\pi/2$ pulse. We then proceed to analyze the QFI under both scenarios, considering the presence and absence of noise.
In Fig. 5a, we observe that without considering any noise channel, spin squeezing yields a higher QFI. However, in the presence of noise (Fig. 5b and Fig. 5c), the advantages of spin squeezing are lost even for small dephasing (Fig. 5b), and Ramsey sequence without spin squeezing is more beneficial. In the case of the qutrit, we observe advantages even in the presence of dephasing (Fig. 2b). Therefore, we can see the benefits of using qutrit sensors over qubit sensors.
@@ V Conclusion
Incorporating a spin squeezing step in magnetometric measurement and utilizing qutrit NV centers have demonstrated a precision increase relative to standard Ramsey magnetometry with qubits and without spin squeezing. However, the advantage of qutrit spin squeezing is notably sensitive to the dephasing environment. Our numerical analysis showed that, as dephasing increases, the advantages become marginal compared to standard methods. We also observed that moderate squeezing results in higher QFI compared to strong squeezing, which is only beneficial for a brief period of free evolution.
Additionally, we have observed that spin squeezing is accompanied by the generation of entanglement, aligning with the general expectation of quantum entanglement as a metrological resource. However, further investigation is needed to determine whether it is spin noise distribution or entanglement alone that causes the precision limit enhancement. Our findings contribute to the broader exploration of quantum technologies beyond qubits, where qutrits and higher-dimensional qubits (qudits) are increasingly considered, particularly in the context of quantum computing [52, 53]. An intriguing question persists regarding the scalability of these advantages, especially regarding the utilization of NV center qutrit ensembles or the investigation of higher-level systems (qudits). These avenues pose exciting challenges for future studies, promising further insights.
Figure 6: The curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.

### v2

@@ I Introduction
In this article, we explore the potential enhancements in quantum metrological performance achievable by transitioning from two-level (qubits) to three-level (qutrits) spin systems and incorporating spin squeezing within typical quantum magnetometry settings, specifically NV center Ramsey interferometry. We adopt a foundational approach, comparing qubit versus qutrit Ramsey quantum magnetometry scenarios, both with and without spin squeezing, in terms of Quantum Fisher Information (QFI). Considering the influence of quantum dephasing and relaxation during the spin system’s evolution, our findings suggest that qutrits, coupled with the additional step of spin squeezing, can outperform qubits in terms of QFI, provided that the spin squeezing magnitude, the timing of squeezing and free evolution are optimally determined.
@@ II Model
In a scenario where there exists an unknown, potentially weak magnetic field within a defined region, our objective is to measure this magnetic field with the utmost precision possible. To accomplish this, we utilize one or two NV centers as our probes. The NV center, characterized as an optically active color defect, results from a substitutional nitrogen impurity and a neighboring carbon vacancy in the diamond lattice. Comprising six electrons from nitrogen and the surrounding carbon atoms, the negatively charged NV center is described as a spin-1 system. We opt to explore the full capabilities of the NV center by modeling it as a qutrit system. This choice is motivated by the potential advantages that qutrits offer in metrology when compared to qubits [29, 30]. Its ground state is a spin triplet (${}^{3}A_{2}$) denoted as |$\ket{Sm_{S}}$ with $S=1$ and $m_{S}=0,\pm 1$. The NV center Hamiltonian captures these properties, while the excited-state triplet (${}^{3}E$) is acknowledged at a higher energy level and is not explored in this context [31, 32].
@@
Figure 1: The application of Ramsey-type sequence on two Nitrogen-Vacancy (NV) Centers with and without the introduction of spin squeezing. $T_{\mathrm{fr}}$ is the free evolution time. $T_{\mathrm{sq}}$ is the duration of the application of spin squeezing.
@@ II.1 Free Hamiltonian and Initial states
When incorporating spin squeezing, the initial state, |$\ket{\psi_{0}}$, is first driven through spin squeezing, and then subjected to the $\pi/2$ pulse, as illustrated in Fig. 1.
@@ II.2 Spin Squeezing Generation
where $c_{1}$ is the squeezing constant. The Hamiltonian, as depicted in Eq. (6), is referred as a one-axis twisting Hamiltonian. The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin. This, combined with the optimal initial state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
@@
Figure 2: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G for the spin-1 formalism. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$.
@@ II.3 Dynamics
The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins. These nuclear spins introduce a dynamic and unpredictable backdrop, contributing to the environmental noise that impacts the precision of measurements performed on the NV center. We assume very low temperatures relative to the NV center transition frequencies. Dephasing and thermalization are incorporated through the dissipators given as follows,
@@ III.2 Spin Squeezing Parameter
with $I_{3}$ being the three-dimensional unit matrix. To calculate spin squeezing, we calculate the mean spin squeezing direction $\hat{n}$ for the state,
@@
And then, we take the minimal variance, $(\Delta\hat{J}^{2}_{n_{\perp}})_{\text{min}}$, in the perpendicular direction of $\hat{n}$. The presence of spin squeezing is characterized by $\xi^{2}<1$. Spin squeezing is closely tied to the creation of entanglement [37, 49, 50]. Additionally, a previous study suggests that initial squeezing can aid in entanglement [51]. Our results are consistent with this statement (Appendix A).
Figure 3: For the spin-1 formalism, the curve in solid blue represents the squeezing parameter, $\xi^{2}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
@@ IV Results
[…]
@@ IV.1 Analysis with spin-1 formalism
Fig. 2 compares one NV center without spin squeezing (blue), two NV centers without spin squeezing (red) and two NV centers with spin squeezing applied at varying strengths, specifically with squeezing constants of $c_{1}=0.25$ GHz (green) and $c_{1}=2.5$ GHz (cyan).
In the case without spin squeezing, Fig. 2a shows a twofold increase in the QFI when transitioning from one NV center to two NV centers. When environmental noise is introduced (Fig. 2b and Fig. 2c), whether the QFI increases or decreases for one NV center, the twofold enhancement persists for two NV centers.
In the case with spin squeezing, in the absence of noise (Fig. 2a), the QFI is larger when spin squeezing is applied with squeezing constant, $c_{1}=0.25$ GHz. It increases further with a higher squeezing constant, $c_{1}=2.5$ GHz. When environmental noise is introduced with $\gamma_{\text{d}}=0.1\gamma_{\text{t}}$ (Fig. 2b), spin squeezing remains beneficial, particularly at a lower squeezing constant, $c_{1}=0.25$ GHz. This suggests that in the presence of noise, a higher squeezing constant does not necessarily lead to higher QFI. Instead, a low squeezing constant ($c_{1}=0.25$ GHz), which is sufficient enough to generate spin squeezing, is more advantageous. However, within the first 4 $\mu$s, a higher squeezing constant, $c_{1}=2.5$ GHz, provides more benefit.
Finally, Fig. 2c shows that the advantage of squeezing is lost when the dephasing rate increases to $\gamma_{\text{d}}=\gamma_{\text{t}}$. While spin squeezing enhance magnetic sensing capabilities, they are also highly sensitive to environmental noise, particularly dephasing. Analyzing the behavior of the QFI curves under such noise conditions is crucial. By identifying the time at which the QFI reaches its maximum, we can optimize experimental design and ensure measurements are performed at the peak sensitivity. Experimentally, we typically measure populations or ($\hat{S}_{z}$) values. However, it is important to note that while the QFI is calculated for the optimal Positive Operator-Valued Measure (POVM), specifically the eigenprojectors of the SLD, measuring $\hat{S}_{z}$ could potentially be suboptimal.
The effect of spin squeezing on the NV centers is further illustrated in Fig. 3. The squeezing parameter, initially around 0.5 (below 1) at the beginning of free evolution, increases due to dephasing and thermalization during this period. Appendix A demonstrates that spin squeezing also leads to entanglement.
Furthermore, given the initial state is oriented along the $z$-axis, when spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis. A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid. The effect of the magnetic field $B$ is to precess the spin around the $z$-axis, and the $x$-axis twisting deforms the noise cone, effectively narrowing the spin’s error ellipse. This sequence is illustrated in Fig. 4.
@@
Figure 5: The Quantum Fisher Information (QFI) as a function of time with $B=50$ G for the spin-1/2 formalism. The curves represents the free evolution as follows: the solid blue curve corresponds to a single NV center, the dashed-dotted red curve corresponds to two NV centers, the pointed cyan curve represents two NV centers subjected to spin squeezing ($c_{1}=2.5$ GHz, $T_{\mathrm{sq}}=3$ ns), and the dashed green curve represents two NV centers subjected to spin squeezing ($c_{1}=0.25$ GHz, $T_{\mathrm{sq}}=30$ ns). The different scenarios are illustrated as: (a) without including any noise terms, (b) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\gamma_{t}$, and (c) including thermalization and dephasing with $\gamma_{t}=0.2$ MHz and $\gamma_{d}=\gamma_{t}$.
@@ IV.2 Analysis with spin-1/2 formalism
As in the spin-1 formalism, if spin squeezing is added, it is applied before the $\pi/2$ pulse. We then proceed to analyze the QFI under both scenarios (with and without squeezing), considering the presence and absence of noise.
In Fig. 5a, we observe that without considering any noise channel, spin squeezing yields a higher QFI. However, in the presence of noise (Fig. 5b and Fig. 5c), the advantages of spin squeezing are lost even for small dephasing (Fig. 5b), and Ramsey sequence without spin squeezing is more beneficial. Earlier in Fig. 2b, for the qutrit case, we observed advantages even in the presence of dephasing. Therefore, the benefits of using qutrit sensors over qubit sensors are clearly evident.
@@ V Conclusion
Incorporating a spin squeezing step in magnetometric measurement and utilizing qutrit NV centers have demonstrated a precision increase relative to standard Ramsey magnetometry. However, the advantage of spin squeezing is notably sensitive to the dephasing environment. Our numerical analysis showed that, as dephasing increases, the advantages become marginal compared to standard methods. We also observed that moderate squeezing results in higher QFI compared to strong squeezing, which is only beneficial for a brief period of free evolution. Furthermore, by conducting the entire analysis using the spin-1/2 formalism and comparing it with the spin-1 formalism, we have demonstrated the advantages of utilizing a qutrit system over a qubit system.
Figure 6: For the spin-1 formalism, the curve in solid blue represents the logarithmic negativity, $E_{N}$, as a function of time during the free evolution of two NV centers subjected to spin squeezing with $c_{1}=2.5$ GHz for a time of $T_{\mathrm{sq}}=3$ ns, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G.
Additionally, we have observed that spin squeezing is accompanied by the generation of entanglement, aligning with the general expectation of quantum entanglement as a metrological resource. However, further investigation is needed to determine whether it is spin noise distribution or entanglement alone that causes the precision limit enhancement. Our findings contribute to the broader exploration of quantum technologies beyond qubits, where qutrits and higher-dimensional qubits (qudits) are increasingly considered, particularly in the context of quantum computing [52, 53]. An intriguing question persists regarding the scalability, as outlook in [54], of these advantages, especially regarding the utilization of NV center qutrit ensembles or the investigation of higher-level systems (qudits). These avenues pose exciting challenges for future studies, promising further insights.
Acknowledgements
This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBITAK) under Project Number 123F150. L. G. and Ö. E. M. thank TÜBITAK for their support.